# Recall, Robustness, and Lexicographic Evaluation

### v5

@@ 6.1.1. Data
We evaluated ranking system runs across a variety of conditions (Table 5). For each dataset, we have a set of evaluation requests and associated relevance judgments. In addition, each dataset involved a number of competing systems, each of which produced a ranking for every request. All datasets were downloaded from NIST, except for movielens, libraryThing, and beerAdvocate, which were downloaded from a public repository. [^8] https://github.com/dvalcarce/evalMetrics For these recommender systems datasets, consistent with (Valcarce et al., 2020), we converted graded judgments to binary labels by considering any rating below 4 as nonrelevant and otherwise relevant for these datasets. In order to analyze results for different retrieval depths, we categorized TREC datasets as deep ($\tilde{n}>1000$), standard ($\tilde{n}\in(100\mathrel{{.}\,{.}}\nobreak 1000]$), or shallow ($\tilde{n}\leq 100$).
Table 5. Datasets used in empirical analysis. Runs submitted to the associated TREC track or recommendation task. TREC tracks are labeled according to the depth of runs: ‘deep’ ($\tilde{n}>1000$), ‘standard’ ($\tilde{n}\in(100\mathrel{{.}\,{.}}\nobreak 1000]$), ‘shallow’ ($\tilde{n}\leq 100$); recommendation tasks are labeled separately as ‘recsys’ for comparison.
@@
legal (2006) 39 34 110.85 4835.07 deep
legal (2007) 43 68 101.023 22240.30 deep
@@
robust (2004) 249 110 69.93 913.82 standard
movielens 6005 21 18.87 100.00 recsys
libraryThing 7227 21 13.15 100.00 recsys
beerAdvocate 17564 21 13.66 99.39 recsys
@@ 6.2.1. Agreement with Existing Metrics
To understand the similarity of lexirecall and traditional metrics, we measured its preference agreement with traditional metrics. Specifically, given an observed a metric difference, $\mu(\pi)\neq\mu(\pi^{\prime})$, for a traditional metric in our datasets, we computed how often lexirecall agreed with the ordering of $\pi$ and $\pi^{\prime}$,
@@
Figure 10. Agreement of system ranking between lexirecall and traditional metrics using Kendall’s $\tau$.
Figure 10 shows the agreement between the ranking of systems by lexirecall and traditional metrics for each dataset, as measured by Kendall’s $\tau$. We aggregated per-query lexirecall system ordering using MC4 (Dwork et al., 2001), as suggested by earlier work in preference-based evaluation (Diaz and Ferraro, 2022; Diaz, 2023). The results are largely consistent with Figure 9. We do not observe perfect correlation with $\text{R}_{1000}$ as we did Figure 9 because 1 scalar $\text{R}_{1000}$ values introduce noise and 2 aggregated lexirecall includes preferences when $\text{R}_{1000}$ is tied. Nevertheless, the agreement is still high relative to other metrics.
@@ 6.2.2. Detection of Difference Between Rankings
Figure 11. Fraction of tied comparisons of rankings $\pi,\pi^{\prime}$ in TREC datasets.
@@ 6.2.3. Statistical Sensitivity
We are also interested in ability of lexirecall to detect statistical differences between pairs of runs (i.e. sets of rankings generated by a single system for a shared set of queries). To do so, we adopted Sakai’s method of measuring the discriminative power of a metric (Sakai, 2014). This approach measures the fraction of pairs of systems that the method detects statistical differences with $p<0.05$. We use two methods to compute $p$-values. In the first, we compute a standard statistical test and, correcting for multiple comparisons, measure the fraction of $p$-values below $0.05$. For lexirecall, we adopt a binomial test since we have binary outcomes. For other metrics, we adopt a Student’s $t$-test, as recommended in the literature (Smucker et al., 2007). We also conducted experiments using incorrect-but-consistent statistical tests with similar outcomes. In order to correct for multiple comparisons for all tests, we use the conservative Holm-Bonferroni method (Boytsov et al., 2013). Our second method of computing $p$-values uses Tukey’s honestly significant difference (HSD) test as proposed by Carterette (2012). This method is considered a more appropriate approach to addressing multiple comparisons compared to our first approach. The goal of this analysis is to understand the statistical sensitivity of lexirecall compared to other recall-oriented metrics, while presenting non-recall metrics for reference.
We present the results of this analysis in Figure 12. When using standard tests (Figure 11(a)), lexirecall is slightly better at detecting significant differences compared to existing recall metrics at deeper retrievals. We can refine this analysis by inspecting the HSD results (Figure 11(b)). In this case, the sensitivity of lexirecall manifests more strongly, clearly more discriminative than existing recall metrics for deep retrievals, although losing this power as retrieval depth decreases. This is consistent with previous observations for preference-based evaluation (Diaz and Ferraro, 2022; Diaz, 2023).
(a) Paired Test
(b) HSD
Figure 12. Statistical sensitivity. Fraction of run pairs where we observe a statistically significant difference (i.e. $p<0.05$).
@@ 6.2.4. Label Degradation
Effective evaluation methods are stable in the presence of missing relevance labels. In this analysis, we held the number of queries fixed and randomly removed a fraction of relevant items per query, leaving at least one relevant item per query. We consider two ways to sample items to remove. First, we uniformly sample from all relevant items $\mathcal{R}$ to remove. Second, we sample from $\mathcal{R}$ based on the number of rankings in $\tilde{S}_{n}$ the item appears in. We expect metrics to degrade in performance more quickly when removing ‘popular’ items. We present results for the effect of label degradation on the fraction of ties (we expect more ties with fewer labels) and preference agreement with full data (we expect lower agreement with fewer labels). As with the previous section, the goal of this analysis is to compare lexirecall to other recall-oriented metrics, while presenting non-recall metrics for reference.
@@
In terms of fraction of ties (Figure 13), lexirecall degrades comparably to metrics like AP and $\mathrm{NDCG}$ and substantially more gracefully compared to existing recall metrics $\text{R}_{1000}$ and RP. While the importance of relevance labels for recall-oriented evaluation is important, this result suggests that existing metrics are extremely brittle when labels are missing. All methods observed more ties at shallower retrieval depths with degradation more pronounced for traditional recall-oriented metrics.
(a) Uniform Label Degradation
(b) Ranking Frequency Label Degradation
@@
Figure 13. Number of ties as labels removed. For each query, we removed a fraction of items from $\mathcal{R}$ and counted the number of tied rankings. Relevant items were either (12(a)) sampled uniformly at random or (12(b)) weighted by their frequency in amongst the submitted rankings $\tilde{S}_{n}$. Average over ten samples.
(a) Uniform Label Degradation
(b) Ranking Frequency Label Degradation
@@
Figure 14. Preference agreement between metric based on degraded labels and complete labels. For each query, we removed a fraction of items from $\mathcal{R}$ and measured the fraction of preferences $\succ_{\mu}$ based on the incomplete relevance judgments that agreed with the preference when using the complete set of relevant items. Relevant items were either (13(a)) sampled uniformly at random or (13(b)) weighted by their frequency in amongst the submitted rankings $\tilde{S}_{n}$. Average over ten samples.
@@
In terms of agreement with preferences based on full data (Figure 14), lexirecall again degrades comparably to AP and $\mathrm{NDCG}$ while RP is much more sensitive across degradation levels. On the other hand, $\text{R}_{1000}$ behaves similar to lexirecall when preserving most labels, but, for drastically sparse labels, the agreement drops. We again see slightly worse degradations with shallower retrieval depths across all metrics. RP in particular demonstrates significantly worse degradation compared to all metrics, while $\text{R}_{1000}$ shows worse degradation when removing relevant items based on ranking frequency.
@@ 7.1.1. Labeling
Although lexirecall appears more stable in the presence of missing labels than existing recall-oriented metrics, the performance of recall and robustness evaluation depends critically on having comprehensive relevance labels. While time-consuming, we believe that, in order to develop robust and fair systems, new techniques for expanding labeled sets for recall evaluation are necessary. Initiatives like TREC adopt pooling as a strategy to achieve more complete judgments. Unfortunately, in situations where relevance is derived from behavioral feedback (e.g. (Kelly and Teevan, 2003; Chandar et al., 2020)), comprehensiveness of relevance is often not the focus. Moreover, the transience of information needs in production environments makes reliable detection of $\mathcal{R}$ an open problem. This situation is exacerbated in recommender system environments where judgments, while often highly personalized and based on psychological relevance, can be extremely incomplete (see (Herlocker et al., 2004) for a discussion).
@@ 7.2.3. A Comment on Graded Metrics
Although we have focused on binary relevance, many ranking scenarios use graded or ordinal relevance. Consider relevance labels represented as an ordinal scale, where higher grades reflect a higher probability of satisfying the information need according to the rater’s subjective opinion. Under such a grading scheme, an item labeled with the minimum grade has a probability of relevance of 00 (i.e. no user would ever find the item relevant) and an item labeled with the maximum grade has a probability of relevance of $1-\epsilon$ (i.e. almost every user would find the item relevant). We can determine grades that reflect the probability of relevance through 1 labeling instructions (e.g. ‘an item with a high grade should satisfy many users; an item with a medium grade should satisfy some users; an item with a low grade should satisfy few users’), 2 voting schemes (Gordon et al., 2021), or 3 aggregated behavioral data (e.g., clickthrough rate) (Zheng et al., 2010). No matter how grades are determined, for a fixed request, a user with a less popular intent may not be satisfied by an item relevant to a more popular intent. This implies that notions of optimality in graded ranking evaluation (i.e. that higher grades should be ranked above lower grades) explicitly values dominant group intents over minority intents. From the perspective of robustness, this means that, for graded judgments, we should consider $\mathcal{R}$ to include all items with a non-zero chance of satisfying a user. This is precisely the approach adopted for $\operatorname{TSE}$ and lexirecall.
@@ 7.2.4. Robustness of Optimal Rankings
We present a proof in Appendix C. Figure 15 displays the difference in worst-case performance between optimal deterministic and stochastic rankings for $m\in[1\mathrel{{.}\,{.}}\nobreak 25]$. This result provides evidence from a robustness perspective that ranking system design should explore the design space of stochastic rankers.
@@ 7.3.1. Recovering an Evaluation Metric
where $\Delta=(n+\epsilon)^{-1}$ and $\epsilon\in(0,1)$ is a free parameter.
We can then define metric lexirecall as,
@@ 8. Conclusion
Appendix A Metric Properties

### v6

@@ 6.1.1. Data
We evaluated ranking system runs across a variety of conditions (Table 5). For each dataset, we have a set of evaluation requests and associated relevance judgments. In addition, each dataset involved a number of competing systems, each of which produced a ranking for every request. The movielens, libraryThing, and beerAdvocate datasets were downloaded from a public repository with splits and processing described in prior work (Valcarce et al., 2020). [^10] https://github.com/dvalcarce/evalMetrics Amazon data sets were prepared using LensKit 0.14.4 (Ekstrand, 2020) to train a variety of recommendation models on 5-core global split datasets from the 2023 Amazon Reviews data (Hou et al., 2024) and generate top-K recommendation runs for the users in the test splits, using an early version of the LensKit Codex [^11] https://codex.lenskit.org; runs will be provided as supplemental resources for this paper when published.. For recommender systems datasets, consistent with (Valcarce et al., 2020), we converted graded judgments to binary labels by considering any rating below 4 as nonrelevant and otherwise relevant for these datasets. All retrieval datasets were downloaded from NIST. In order to analyze results for different ranking depths, we categorized datasets as deep ($\tilde{n}>1000$), standard ($\tilde{n}\in(100\mathrel{{.}\,{.}}\nobreak 1000]$), or shallow ($\tilde{n}\leq 100$).
Table 5. Datasets used in empirical analysis. Runs submitted to the associated TREC track or recommendation task. Datasets are labeled according to the depth of runs: ‘deep’ ($\tilde{n}>1000$), ‘standard’ ($\tilde{n}\in(100\mathrel{{.}\,{.}}\nobreak 1000]$), ‘shallow’ ($\tilde{n}\leq 100$).
@@
recommendation
amzn-cds-vinyl-100 6875 10 2.76 100 shallow
amzn-cds-vinyl-1000 6875 10 2.76 1000 standard
amzn-cds-vinyl-2000 6875 10 2.76 2000 deep
amzn-musical-instruments-100 11648 10 3.08 100 shallow
amzn-musical-instruments-1000 11648 10 3.08 1000 standard
amzn-musical-instruments-2000 11648 10 3.08 2000 deep
amzn-software-100 6345 10 2.28 100 shallow
amzn-software-1000 6345 10 2.28 1000 standard
amzn-software-2000 6345 10 2.28 2000 deep
amzn-video-games-100 11554 10 3.08 100 shallow
amzn-video-games-1000 11554 10 3.08 1000 standard
amzn-video-games-2000 11554 10 3.08 2000 deep
beerAdvocate 17564 21 13.66 99.39 shallow
libraryThing 7227 21 13.15 100.00 shallow
movielens 6005 21 18.87 100.00 shallow
retrieval
@@
deep-docs (2022) 76 42 1245.62 98.86 shallow
deep-docs (2023) 82 5 75.10 100.00 shallow
@@
deep-pass (2022) 76 100 628.145 97.50 shallow
deep-pass (2023) 82 35 49.87 99.90 shallow
legal (2006) 39 34 110.85 4835.07 deep
legal (2007) 43 68 101.023 22240.30 deep
robust (2004) 249 110 69.93 913.82 standard
@@ 6.2.1. Agreement with Existing Metrics
To understand the similarity of lexirecall and traditional metrics, we measured its preference agreement with traditional metrics. Specifically, given an observed metric difference, $\mu(\pi)\neq\mu(\pi^{\prime})$, for a traditional metric in our datasets, we computed how often lexirecall agreed with the ordering of $\pi$ and $\pi^{\prime}$,
@@ 6.2.2. Detection of Difference Between Rankings
Although the general trends are consistent across recommendation and retrieval contexts, recommendation metrics have substantially higher variance. This results from the sparser set of labels in recommendation, where the total number of unique vectors, ($n\choose m$, is lower and therefore there is a higher probability of ties.
(a) Recommendation
(b) Retrieval
Figure 11. Fraction of tied comparisons of rankings $\pi,\pi^{\prime}$ in recommendation and retrieval datasets.
@@ 6.2.3. Statistical Sensitivity
We are also interested in the ability of lexirecall to detect statistical differences between pairs of runs (i.e. sets of rankings generated by a single system for a shared set of queries). To do so, we adopted Sakai’s method of measuring the discriminative power of a metric (Sakai, 2014). This approach measures the fraction of pairs of systems that the method detects statistical differences with $p<0.05$. We use two methods to compute $p$-values. In the first, we compute a standard statistical test and, correcting for multiple comparisons, measure the fraction of $p$-values below $0.05$. For lexirecall, we adopt a binomial test since we have binary outcomes. For other metrics, we adopt a Student’s $t$-test, as recommended in the literature (Smucker et al., 2007). We also conducted experiments using incorrect-but-consistent statistical tests with similar outcomes. In order to correct for multiple comparisons for all tests, we use the conservative Holm-Bonferroni method (Boytsov et al., 2013). Our second method of computing $p$-values uses Tukey’s honestly significant difference (HSD) test as proposed by Carterette (2012). This method is considered a more appropriate approach to addressing multiple comparisons compared to our first approach. The goal of this analysis is to understand the statistical sensitivity of lexirecall compared to other recall-oriented metrics, while presenting non-recall metrics for reference.
We present the results of this analysis in Figures 12 and 13. When using standard tests (Figure 12), lexirecall is slightly better at detecting significant differences compared to existing recall metrics at deeper retrievals. We can refine this analysis by inspecting the HSD results (Figure 13). In this case, the sensitivity of lexirecall manifests more strongly, clearly more discriminative than existing recall metrics for deep retrievals, although losing this power as retrieval depth decreases. This is consistent with previous observations for preference-based evaluation (Diaz and Ferraro, 2022; Diaz, 2023).
(a) Recommendation
(b) Retrieval
Figure 12. Statistical sensitivity. Fraction of run pairs where we observe a statistically significant difference (i.e. $p<0.05$) using a paired test.
(a) Recommendation
(b) Retrieval
Figure 13. Statistical sensitivity. Fraction of run pairs where we observe a statistically significant difference (i.e. $p<0.05$) using Tukey’s honestly significant difference test.
@@ 6.2.4. Label Degradation
Effective evaluation methods are stable in the presence of missing relevance labels. In this analysis, we held the number of queries fixed and randomly removed a fraction of relevant items per query, leaving at least one relevant item per query. We uniformly sample from all relevant items $\mathcal{R}$ to remove. We present results for the effect of label degradation on the fraction of ties (we expect more ties with fewer labels) and preference agreement with full data (we expect lower agreement with fewer labels). As with the previous section, the goal of this analysis is to compare lexirecall to other recall-oriented metrics, while presenting non-recall metrics for reference.
@@
In terms of fraction of ties (Figure 14), lexirecall degrades comparably to metrics like AP and $\mathrm{NDCG}$ and substantially more gracefully compared to existing recall metrics $\text{R}_{1000}$ and RP. While the importance of relevance labels for recall-oriented evaluation is important, this result suggests that existing metrics are extremely brittle when labels are missing. All methods observed more ties at shallower retrieval depths with degradation more pronounced for traditional recall-oriented metrics. As with our analysis of the number of ties (Section 6.2.2), because of the sparsity of labels in recommendation, we observe much higher variance compared to retrieval, where labels are more complete due to pooling (Spärck Jones and van Rijsbergen, 1975; Zobel, 1998). As such, the recommendation tasks can be seen as operating in the far right-hand portion of retrieval experiments.
(a) Recommendation
(b) Retrieval
@@
Figure 14. Number of ties as labels removed. For each query, we removed a fraction of items from $\mathcal{R}$ and counted the number of tied rankings. Relevant items were sampled uniformly at random. Average over ten samples.
(a) Recommendation
(b) Retrieval
@@
Figure 15. Preference agreement between metric based on degraded labels and complete labels. For each query, we removed a fraction of items from $\mathcal{R}$ and measured the fraction of preferences $\succ_{\mu}$ based on the incomplete relevance judgments that agreed with the preference when using the complete set of relevant items. Relevant items were sampled uniformly at random. Average over ten samples.
@@
In terms of agreement with preferences based on full data (Figure 15), lexirecall again degrades comparably to AP and $\mathrm{NDCG}$. In recommendation experiments, RP and $\text{R}_{1000}$ behave comparably or better than lexirecall, AP, and $\mathrm{NDCG}$. However, this may be an artifact of the poor behavior of RP and $\text{R}_{1000}$ under extremely sparse labels, as demonstrated in earlier experiments. That is, since these metrics observe so many ties and ties increase with label degradation, accurately predicting these preferences under ‘full data’ will be easier with label degradation. We contrast this with retrieval experiments, where labels are more complete. In this condition, RP is much more sensitive to degradation, dropping in performance quickly. On the other hand, $\text{R}_{1000}$ behaves similar to lexirecall when preserving most labels, but, for drastically sparse labels, the agreement drops. We again see slightly worse degradations with shallower retrieval depths across all metrics. RP in particular demonstrates significantly worse degradation compared to all metrics, while $\text{R}_{1000}$ shows worse degradation when removing relevant items based on ranking frequency.
@@ 7.1.1. Labeling
Although lexirecall appears more stable in the presence of missing labels than existing recall-oriented metrics, the performance of recall and robustness evaluation depends critically on having comprehensive relevance labels. This situation is exacerbated in recommender system environments where judgments, while often highly personalized and based on psychological relevance, can be extremely incomplete (see (Herlocker et al., 2004) for a discussion). This suggests two possibilities. Sparsity in recommendation can arise from 1 very niche relevance or 2 severe under-labeling of relevant items. In the first situation, our experiments suggest that lexirecall is substantially more sensitive than existing recall metrics. However, in the second situation is accurate, we can examine metric behavior under retrieval scenarios, where labels are more complete. Indeed, our experiments would suggest that using existing recall metrics for sparse recommendation tasks does not accurately reflect recall under more complete data. This echos similar observations in the recommender system community for addressing data sparsity issues (Ekstrand and Mahant, 2017; Tian and Ekstrand, 2020; Ihemelandu and Ekstrand, 2023).
While time-consuming, we believe that, in order to develop robust and fair systems, new techniques for expanding labeled sets for recall evaluation are necessary. In retrieval contexts, initiatives like TREC adopt pooling as a strategy to achieve more complete judgments (Spärck Jones and van Rijsbergen, 1975). Unfortunately, in situations where relevance is derived from behavioral feedback (e.g. (Kelly and Teevan, 2003; Chandar et al., 2020)), comprehensiveness of relevance is often not the focus. Moreover, the transience of information needs in production environments makes reliable detection of $\mathcal{R}$ an open problem. This is compounded by the desire—in both retrieval and recommendation contexts—to estimate relevance for evaluation and optimization transparently to the user. As such, new methods for estimating relevance of a broader set of items without impacting the user are necessary.
@@ 7.2.3. A Comment on Graded Metrics
Although we have focused on binary relevance, many ranking scenarios use graded or ordinal relevance. Consider relevance labels represented as an ordinal scale, where higher grades reflect a higher probability of satisfying the information need according to the rater’s subjective opinion, as happens in many retrieval scenarios. Under such a grading scheme, an item labeled with the minimum grade has a probability of relevance of 00 (i.e. no user would ever find the item relevant) and an item labeled with the maximum grade has a probability of relevance of $1-\epsilon$ (i.e. almost every user would find the item relevant). We can determine grades that reflect the probability of relevance through 1 labeling instructions (e.g. ‘an item with a high grade should satisfy many users; an item with a medium grade should satisfy some users; an item with a low grade should satisfy few users’), 2 voting schemes (Gordon et al., 2021), or 3 aggregated behavioral data (e.g., clickthrough rate) (Zheng et al., 2010). No matter how grades are determined, for a fixed request, a user with a less popular intent may not be satisfied by an item relevant to a more popular intent. This implies that notions of optimality in graded ranking evaluation (i.e. that higher grades should be ranked above lower grades) explicitly values dominant group intents over minority intents. From the perspective of robustness, this means that, for graded judgments, we should consider $\mathcal{R}$ to include all items with a non-zero chance of satisfying a user. This is precisely the approach adopted for $\operatorname{TSE}$ and lexirecall.
@@ 7.2.4. Robustness of Optimal Rankings
We present a proof in Appendix D. Figure 16 displays the difference in worst-case performance between optimal deterministic and stochastic rankings for $m\in[1\mathrel{{.}\,{.}}\nobreak 25]$. This result provides evidence from a robustness perspective that ranking system design should explore the design space of stochastic rankers.
@@ 7.3.1. Recovering an Evaluation Metric
where $\Delta=(n+\epsilon)^{-1}$ and $\epsilon\in(0,1)$ is a free parameter. We can then define metric lexirecall as,
@@ 8. Conclusion
Appendix A Full Papers Published at the ACM Conference on Recommender Systems that Measure Recall
We downloaded and reviewed all full papers published at the ACM Conference on Recommender Systems (2007-2024), omitting reproducibility, late-breaking results, short, demo, doctoral, or keynote papers. A paper was identified if recall was quantitatively measured as part of experimentation. Conceptual discussions of recall (e.g., in comparison to experimentation metrics) were not considered to use recall. We identified only those papers that specifically name recall (e.g., $\text{R}_{k}$), although some operationalizations of hit rate may be equivalent to recall (Tamm et al., 2021). We present the full papers identified in Table 7.
Table 7. DOIs of full papers published at the ACM Conference on Recommender Systems (2007-2024) that measure recall in experiments.
year total recall DOIs
2007 16 3 10.1145/1297231.12972{48, 50, 51}
2008 20 3 10.1145/1454008.14540{17, 32, 34}
2009 25 4 10.1145/1639714.16397{24, 26, 37, 45}
2010 25 8 10.1145/1864708.18647{21, 28, 31, 32, 41, 44, 45, 47}
2011 22 5 10.1145/2043932.20439{41, 45, 47, 57, 65}
2012 24 11 10.1145/2365952.23659{62, 63, 67, 68, 69, 72, 73, 76, 79, 82, 89}
2013 32 11 10.1145/2507157.2507{160, 170, 171, 172, 182, 184, 186, 209, 210, 211, 215}
2014 35 8 10.1145/2645710.26457{21, 23, 29, 34, 38, 40, 43, 46}
2015 28 7 10.1145/2792838.28001{70, 72, 74, 76, 85, 89, 95}
2016 29 12 10.1145/2959100.29591{33, 37, 46, 47, 49, 51, 57, 67, 70, 78, 80, 82}
2017 26 6 10.1145/3109859.3109{877, 879, 887, 896, 900, 903}
2018 32 11 10.1145/3240323.3240{343, 347, 349, 355, 363, 371, 372, 374, 381, 391, 405}
2019 36 11 10.1145/3298689.334{6996, 7002, 7007, 7009, 7012, 7013, 7026, 7036, 7044, 7058, 7065}
2020 39 12 10.1145/3383313.341{1476, 2232, 2235, 2243, 2247, 2248, 2249, 2258, 2259, 2262, 2265, 2268}
2021 49 19 10.1145/3460231.3474{228, 230, 234, 238, 240, 242, 249, 252, 255, 257, 260, 263, 265, 266, 268, 270, 272, 273, 275}
2022 39 13 10.1145/3523227.3546{752, 754, 755, 760, 762, 763, 768, 770, 771, 775, 782, 784, 785}
2023 47 21 10.1145/3604915.3608{766, 771, 773, 781, 783, 784, 785, 786, 803, 804, 806, 809, 810, 811, 812, 815, 863, 868, 871, 878, 882}
2024 58 24 10.1145/3640457.3688{096, 098, 100, 104, 108, 109, 113, 117, 121, 122, 123, 124, 125, 127, 133, 137, 138, 139, 145, 146, 148, 149, 151, 153}
Appendix B Metric Properties